a 73 kg person expends 400 W when walking on a level treadmill at a speed of 7.2 km/h. When the treadmill is inclined without changing the speed, the person's expended power increases to 600 W. What is the angle of incline of the treadmill by assuming that all the increased output goes into overcoming the force of gravity.

1
Expert's answer
2022-04-11T12:03:31-0400

"F=\frac{W_2-W_1}v,"

"mg\sin\alpha=\frac{W_2-W_1}v,"

"\alpha=\arcsin\frac{W_2-W_1}{mgv}=8.4°."